find g(x+1) for g(x)=x^2 - 2x
Answer by Prithwis(166)

Find g(x+1) for g(x)=x^2-2x
g(x+1)=(x+1)^2-2(x+1)
=> g(x+1)=x^2+2x+1-2x-2
=> g(x+1)=x^2-1
=> g(x+1)=(x+1)(x-1)
So, your answer is g(x+1)=(x+1)(x-1) for when g(x)=x^2-2x